OpenStudy (anonymous):

Given the function k(x) = x2, compare and contrast how the application of a constant, c, affects the graph. The application of the constant must be discussed in the following manners: k(cx) c • k(x)

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

btw i know that k(cx) controls how wide or narrow the graph is. is that correct?

OpenStudy (ranga):

yes.

OpenStudy (anonymous):

ok so how about the second function?

OpenStudy (anonymous):

thats where i got stuck

OpenStudy (ranga):

k(x) = x^2 (It is a parabola. It opens upwards) First one: k(cx) = (c^2)x^2 (This is also a parabola. c^2 is always positive whether c is negative or positive and therefore this parabola also opens upwards. The value of c^2 will determine how narrow or wide the parabola will be. discus the case of |c| < 1 and |c| > 1

OpenStudy (ranga):

k(x) = x^2 (It is a parabola. It opens upwards) Second case: c * k(x) = cx^2. cx^2 is also a parabola. If c is negative, the parabola will open downward. If c is positive the parabola will open upward. Discuss two possibilities: |c| < 1 and |c| > 1

OpenStudy (anonymous):

ohh i see

OpenStudy (anonymous):

so for the first it controls how WIDE the parabola will be. the second one controls whether the parabola will open UPWARD or DOWNWARD

OpenStudy (ranga):

In the first case, k(cx), will compress or stretch the graph horizontally. If |c| < 1 (that is, a fraction) the graph will be stretched horizontally. If |c| > 1 the graph will be compressed horizontally.

OpenStudy (anonymous):

could you sketch if please?

OpenStudy (anonymous):

it**

OpenStudy (anonymous):